3. A concert is being held in your hometown. 1500 tickets are sold. Adult tickets are $10.50 and child’s tickets are $7.50. If a total of $13950 worth of tickets were sold, determine how many adult tickets and how many child tickets were sold.
Answer: Let x be the number of adult tickets and y be the number of child tickets. System of equations:
1st equation:
2nd equation:
Solution: x = and y =

Answer by Boreal(15235)

x+y=1500
10.50x+7.50y=13950
multiply the first one by -7.5 to eliminate y
-7.5x-7.5y=-11250
3x=2700
x=900 adults for $9450
y=600 children for $4500